# Geodesic Slice Sampling on the Sphere

### v3

@@ 4.1 Bingham Distribution
An efficient MCMC algorithm should explore the sphere rapidly by making large leaps from one sample to the next. Again, we see in Fig. 6 a superior performance of geoSSS. The RWMH algorithm achieves only small jumps. As expected, HMC moves more rapidly over the sphere compared to RWMH, but it still cannot compete with the geodesic slice samplers.
We also run similar tests on a more challenging Bingham target with $d=50$ and $\kappa_{d}=300$, i.e., the dimension of the sample space is much larger and the distribution is more concentrated. We observe the same trends as before. Supplementary Figure A1 shows the distribution of samples projected onto the first mode. Now, both RWMH and HMC are stuck in the first mode and fail to find the second mode, whereas the geodesic slice samplers represent both modes accurately. The ACF of HMC outperforms RWMH as expected (see Supplementary Fig. A2), but still the geodesic slice samplers show a faster decorrelation than HMC resulting in a higher effective sample size. Since no jumps occur during the entire run of RWMH and HMC, their hopping frequencies are estimated to be smaller than $10^{-5}$, whereas geoSSS still achieves an acceptable jump rate (see Appendix Fig. A2). The rejection-based geoSSS clearly outperforms the shrinkage strategy on this target by achieving a higher hopping rate and as a consequence also a more favorable distribution of step sizes $\delta(x_{n+1},x_{n})$ (see Appendix Fig. A3).
4.2 Mixture of von Mises-Fisher Distributions
We test the slice samplers as well as RWMH and HMC also on a $K$-component mixture model of von Mises-Fisher (vMF) distributions in $d$ dimensions. The vMF distribution is defined by the unnormalized density
$p_{\text{vMF}}{}(x;\kappa,\mu)=\exp(\kappa\mu^{T}\!x),\quad x\in\mathbb{S}^{d-1},$ (14)
where $\kappa>0$ is the concentration parameter and $\mu\in\mathbb{S}^{d-1}$. For $K\in\mathbb{N}$ our target distribution is a particular mixture of vMF distributions where each component has the same weight $1/K$ and concentration parameter $\kappa$, i.e., the corresponding unnormalized density takes the form
$p_{\text{mix}}(x)=\frac{1}{K}\sum_{k=1}^{K}p_{\text{vMF}}(x;\kappa,\mu_{k}),\quad x\in\mathbb{S}^{d-1},$ (15)
where every $\mu_{k}\in\mathbb{S}^{d-1}$ with $k\in\{1,\dots,K\}$ is sampled w.r.t. the uniform distribution on the unit sphere and then fixed. To generate a baseline, we use the method by Wood (1994).
As before, we evaluate both geoSSS variants against RWMH and HMC on a $10$-dimensional mixture of vMF distributions with $K=5$ mixture components on the unit sphere $\mathbb{S}^{9}$. We set $\kappa=100$ and draw $10^{6}$ samples.
As a measure for the quality of approximate sampling we consider the distribution with which the samplers visit the $K$ modes of the target mixture model (15). Ideally, this distribution should be uniform, because each component of the mixture has the same weight and concentration parameter. We contrast the empirical frequency $q_{k}$ with which the $k$-th mode is visited by a sampler with the uniform distribution by using the Kullback-Leibler (KL) divergence
$\text{KL}(q\mid p):=\sum_{k=1}^{K}q_{k}\log(q_{k}/p_{k})$
where $p_{k}=1/K$ is the ideal distribution. To estimate $q_{k}$, we assign samples $x_{n}$ to the nearest of the $K$ peaks based on the geodesic distance. This allows us to count how often each mode was visited by the Markov chain. A good MCMC sampler should produce $q_{k}\approx 1/K$ resulting in $\text{KL}(q\mid p)\approx 0$. The more the KL divergence differs from the minimum value of zero, the greater is the mismatch between the ideal and empirical frequency of mode visits. Remarkably as seen from the right panel of Fig. 7, only the geoSSS variants achieve a small KL divergence. RWMH and HMC produce large KL divergences, indicating a poorer representation of the target distribution compared to geoSSS. Although RWMH is the fastest to decorrelate (see left panel of Fig. 7), it is unreliable as observed from the KL divergence result. The geoSSS variants decorrelate much faster as compared to HMC. This aligns with the observations from the marginal histograms (see Fig. 8). Both variants of geoSSS approximates the underlying mixture significantly better compared to HMC and especially RWMH which fails to approximate every component correctly. Furthermore, the great circle distance between successive samples (see Fig. 9) shows that the slice samplers explore the entire sphere again consistent with previous numerical tests.
Figure 7: Left panel: ACF plotted for the first dimension from the $10$-dimensional mixture of a vMF distribution. Right panel: KL divergence between the empirical and ideal frequency with which each mode is visited by the four MCMC samplers.
Figure 8: Marginal distributions (shown in grey) for the mixture of vMF distribution ($d=10,\kappa=100,K=5$) where columns corresponds to MCMC methods and rows correspond to the dimension of the marginal. The red dashed line indicates the baseline obtained with the acceptance/rejection sampler (see Wood, 1994).
Figure 9: Geodesic distance (log-scale) between successive approximate samples for the mixture of vMF distribution ($d=10$, $K=5$ and $\kappa=100$).
@@ 4.2 Mixture of von Mises-Fisher Distributions
We further evaluate the performance of the samplers by fixing the components on the sphere and increasing the concentration parameter $\kappa$ from $50$ to $500$, therefore making the distributions “spikier”. We estimate the effective sample size (ESS) for each method by considering $10$ chains and $10^{6}$ MCMC steps from the first dimension, resulting in ESS values per method for each $\kappa$. Left panel of Fig. 10 shows the estimated ESS for each of the four samplers as a function of the concentration parameter $\kappa$. Overall, we observe that sampling the mixture model becomes more challenging as the components become more concentrated with increasing $\kappa$ for all methods. However, both geoSSS outperform RWMH and HMC, particularly for lower $\kappa$ values. At large $\kappa$ values, all samplers perform similarly poorly and fail to produce a reliable approximation of the target.
@@
Figure 10: Comparing the MCMC samplers for the mixture of vMF distribution ($d=10$, $K=5$) by varying $\kappa$ from 50 to 500; Left panel: Considering the first dimension, estimating the relative effective sample size (ESS) for 10 chains. Right panel: Number of rejections per MCMC step.
@@
It is worth noting that the rejection-based geoSSS yields the highest ESS, albeit at the cost of generating a significantly larger number of rejected samples in comparison to the shrinkage-based geoSSS. This is demonstrated in the right panel of Fig. 10, where it is observed that the rejection-based geoSSS generates an increasingly larger number of rejected samples for higher values of $\kappa$. For instance, for $\kappa=50$, approximately 17 rejections happen per MCMC step, while for $\kappa=500$, around 60 rejections occur. On the other hand, the number of rejections for the shrinkage-based geoSSS increases at a much slower rate with $\kappa$. It can be seen here that for $\kappa=50$, we have approximately 4 rejections per MCMC step, and for $\kappa=500$, approximately just 6 rejections per MCMC step occur. Thus, selecting between the two samplers necessitates a trade-off between computational requirements and accuracy. The number of rejections directly determines the number of log probability evaluations and thereby the computational costs of the slice samplers. RWMH and HMC, on the other hand, have a fixed computational budget per MCMC step. In case of RWMH, one step requires a single log probability evaluation, whereas an HMC step involves one log probability evaluation as well as multiple gradient evaluations, one for each leapfrog step. In our settings, we use 10 leapfrog integration steps. Among the four MCMC strategies studied here, RWMH has the lowest computational costs, but also the smallest ESS. The computational costs of shrinkage-based geoSSS is smaller than the costs of HMC, nevertheless geoSSS also performs better in terms of the other evaluation criteria in comparison to HMC.
4.3 Curved Distribution on the 2-Sphere
@@
HMC is expected to show superior performance on target densities that accumulate ‘mass’ along narrow connected regions. To study such a sampling problem, we define a spherical distribution along a curve. The curve is created by picking 10 points uniformly distributed on the $2$-sphere and connecting them such that the overall distance is shortest (this is done by solving a small traveling salesman problem). We use spherical linear interpolation (slerp) to interpolate between two successive points $x,y\in\mathbb{S}^{d-1}$, i.e., we apply
@@
where $\theta=\arccos(x^{T}y)$, cf. Hanson (1995). If $\mu(t)$ with $t\in[0,1]$ is the curve obtained by concatenating the slerps between the 9 pairs of successive points, then we define the spherical distribution with unnormalized density
$p(x)=\exp\left\{\kappa\max_{t\in[0,1]}x^{T}\!\mu(t)\right\},\qquad x\in\mathbb{S}^{d-1},$ (16)
which we also call curved von Mises-Fisher distribution (curved vMF).
@@
The maximum of $x^{T}\!\mu(t)$ for a single slerp can be computed in closed form. The total maximum over the entire curve is the maximum over the maxima of slerps connecting two successive points. The probability distribution determined by (16) concentrates ‘mass’ around the curve. As in the case for the standard vMF distribution, the parameter $\kappa$ controls the concentration. In our sampling tests, we choose $\kappa=300$. We run all four samplers on targeting the corresponding curved vMF. Rejection-based geoSSS and HMC takes roughly similar run duration, HMC taking slightly longer. However, almost twice as long as shrinkage-based geoSSS. RWMH is by far the fastest method. Refer Appendix J for run times.
Figure 11: Samples of the four different MCMC samplers targeting a curved vMF on the $2$-sphere are shown as black dots. The red curve is generated by positioning $10$ points uniformly distributed on the sphere and interpolating between them using slerp. The concentration parameter is $\kappa=300$.
Figure 11 shows the samples of the four Markov chain approaches of the first $10000$ MCMC iterations. RWMH fails to explore the entire curve. On the other hand, the slice samplers and, as expected, HMC manage to explore the entire curve fairly uniformly. To measure the discrepancy between the discretized target and the histogram generated by a sampling procedure, we use KL divergence as a metric. Right panel of Fig. 12 shows that the two slice samplers achieve the lowest KL values, surprisingly better than HMC, and thus produce a representation of the target that is more faithful to the true distribution than what RWMH is able to achieve. RWMH also shows the worst ACF with a very slow decay (left panel Fig. 12). Among the other three samplers, the slice samplers outperform HMC in that they decorrelate more rapidly. As expected, the rejection-based geoSSS shows the fastest decorrelation. The distribution of distances between successive samples also disfavors RWMH (Fig. 13). Both geoSSS variants show a peak at very short distances. HMC performs better in that it avoids minor changes and produces larger distances. Nevertheless the slice samplers reach out to larger distances covering the entire range of great-circle distance from 0 to $\uppi$ more often than HMC.
@@
Figure 12: Left panel: ACF plotted for the first dimension based on $10^{5}$ samples. Right panel: KL divergence between a discretized version of the curved vMF target and a histogram compiled from the MCMC samples.
Figure 13: Geodesic distance (log-scale) between successive approximate samples for the curved von Mises-Fisher distribution (16).
@@ 5 Summary
We introduce two slice sampling based MCMC-methods on the sphere, the ideal geodesic and geodesic shrinkage slice samplers, that use movements on great circles. For the Markov kernels of both samplers we are able to establish reversibility and explicit convergence statements under mild assumptions. In numerical experiments we see that in particular for moderately concentrated target distributions with a moderate number of dimensions our slice samplers perform well and can compete with RWMH and HMC on the sphere, while having the additional advantage of being free of tuning parameters. For multimodal target distributions we observe that the ideal geodesic slice sampler and the geodesic shrinkage slice sampler even outperform RWMH and HMC. However, we observe that the performance of our methods, just as for RWMH and HMC, deteriorates for increasing concentration of the target distribution, and it is to be expected that the same happens for increasing dimension. This dependence on concentration and dimension already appears in Theorem 6 and Theorem 16. Regarding the comparison between the two geodesic slice samplers, the ideal version seems to outperform the shrinkage based sampler in terms of Markov chain transitions. However, this comes with a (significantly) greater computational cost per transition.
@@
The implementation of our geoSSS algorithms as well as RWMH (see Appendix F) and spherical HMC (Lan et al., 2014) on the sphere as well as scripts for all the numerical illustrations are compiled in our package labeled “GeoSSS”, accessible at this GitHub repository https://github.com/microscopic-image-analysis/geosss. The numerical illustration results computed from the “/scripts” directory from the repository are uploaded at Zenodo https://doi.org/10.5281/zenodo.8287302.
@@ Appendix A Simulation of $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$
Implementing the slice sampling algorithms proposed in this paper involves sampling from $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}=\frac{1}{\omega_{d-2}}\,\mu_{x}$ for $x\in\mathbb{S}^{d-1}$. A scheme for performing it is provided in Algorithm 5 and justified as follows. It exploits that ${\mathbb{S}^{d-2}_{x}}$ can be interpreted “as tilted $\mathbb{S}^{d-1}$, such that it lies in the $(d-1)$-dimensional hyperplane orthogonal to $x$”. We formalize the arguments and require further objects. For $x\in\mathbb{S}^{d-1}$ let
@@
be the $(d-1)$-dimensional hyperplane with normal vector $x$. Since this is a $(d-1)$-dimensional subvectorspace of $\mathbb{R}^{d}$, we can choose an orthonormal basis $e_{1}^{(x)},\ldots,e_{d-1}^{(x)}\in E_{x}$ of $E_{x}$. The isometric mapping
@@
describes the rotation of $\mathbb{R}^{d-1}$ into $E_{x}$. Therefore it assigns the unit sphere in $\mathbb{R}^{d-1}$ to the unit sphere in $E_{x}$, i.e., $\varphi_{x}(\mathbb{S}^{d-2})={\mathbb{S}^{d-2}_{x}}$. Isometries preserve volume, such that
@@
We prove the validity of Algorithm 5 based upon this relationship between $\sigma_{d-2}$ and $\mu_{x}$. More detailed, for $x\in\mathbb{S}^{d-1}$ we establish an expression for the mapping $y\mapsto y-x^{T}yx$ in terms of $\varphi_{x}$ that describes the projection of $\mathbb{R}^{d}$ onto $E_{x}$. Then, we deduce that Algorithm 5 realizes a transformed normalized standard normally distributed random variable under the isometry $\varphi_{x}$. By (18) this implies the desired result formalized as follows.
@@
$y-{x}^{T}\!{y}\,x$ $=\sum_{i=1}^{d-1}\alpha_{i}e_{i}^{(x)}=\varphi_{x}\left((\alpha_{1},\dots,\alpha_{d-1})^{T}\right)=\varphi_{x}(U^{T}y).$

### v4

@@ 4.2 Mixture of von Mises-Fisher Distributions
An efficient MCMC algorithm should explore the sphere rapidly by making large leaps from one sample to the next. The great circle distance between successive samples (see Fig. 6C) shows that the slice samplers explore the sphere more efficiently than HMC and RWMH.
@@
We further evaluate the performance of the samplers by fixing the components on the sphere and increasing the concentration parameter $\kappa$ from $50$ to $500$, therefore making the distributions “spikier”. We estimate the effective sample size (ESS) for each method by considering $10$ chains and $10^{6}$ MCMC steps from the first dimension, resulting in ESS values per method for each $\kappa$. Figure 7A shows the estimated ESS for each of the four samplers as a function of the concentration parameter $\kappa$. Overall, we observe that sampling the mixture model becomes more challenging as the components become more concentrated with increasing $\kappa$ for all methods. However, both geoSSS outperform RWMH and HMC, particularly for lower $\kappa$ values. At large $\kappa$ values, all samplers perform similarly poorly and fail to produce a reliable approximation of the target.
@@
Figure 7: Comparing the MCMC samplers for the mixture of vMF distribution ($d=10$, $K=5$) by varying $\kappa$ from 50 to 500. (A) Considering the first dimension, estimating the relative effective sample size (ESS) for 10 chains. (B) Number of rejections per MCMC step.
@@
It is worth noting that the rejection-based geoSSS yields the highest ESS, albeit at the cost of generating a significantly larger number of rejected samples in comparison to the shrinkage-based geoSSS. This is demonstrated in Fig. 7B, where it is observed that the rejection-based geoSSS generates an increasingly larger number of rejected samples for higher values of $\kappa$. For instance, for $\kappa=50$, approximately 17 rejections happen per MCMC step, while for $\kappa=500$, around 60 rejections occur. On the other hand, the number of rejections for the shrinkage-based geoSSS increases at a much slower rate with $\kappa$. It can be seen here that for $\kappa=50$, we have approximately 4 rejections per MCMC step, and for $\kappa=500$, approximately just 6 rejections per MCMC step occur. The number of rejections directly determines the number of log probability evaluations and thereby the computational costs of the slice samplers. RWMH and HMC, on the other hand, have a fixed computational budget per MCMC step. In case of RWMH, one step requires a single log probability evaluation, whereas an HMC step involves one log probability evaluation as well as multiple gradient evaluations, one for each leapfrog step. In our settings, we use 10 leapfrog integration steps. In summary, RWMH has the lowest computational costs, but also the smallest ESS. The computational costs of shrinkage-based geoSSS is smaller than the costs of HMC, whilst it also performs better than HMC in terms of the other evaluation criteria.
4.3 Curved Distribution on the Sphere
@@
For a target density where the ‘mass’ accumulates along narrow connected regions, HMC is expected to show superior performance. To study such a sampling problem, we define a spherical distribution along a curve. The curve is created by picking 10 consecutive points on the (d1)(d-1)-sphere and to ensure a smooth path, we connect every two successive points $x,y\in\mathbb{S}^{d-1}$ using spherical linear interpolation (slerp), i.e., we apply
@@
as map from $[0,1]$ to $\mathbb{S}^{d-1}$, where $\theta=\arccos(x^{T}y)$, cf. (Hanson, 1995). If $\mu(t)$ with $t\in[0,1]$ is the curve obtained by concatenating and rescaling the slerps between the 9 pairs of successive points, then we define the spherical distribution with unnormalized density
$p(x)=\exp\left(\kappa\max_{t\in[0,1]}x^{T}\!\mu(t)\right),\qquad x\in\mathbb{S}^{d-1},$ (17)
@@
which we also call curved von Mises-Fisher distribution (curved vMF). The maximum of $x^{T}\!\mu(t)$ for a single slerp can be computed in closed form. The total maximum over the entire curve is then the maximum over the maxima of slerps connecting two successive points. The probability distribution corresponding to (17) concentrates probability mass around the curve $\mu(t)$. As in the case for the standard vMF distribution, the parameter $\kappa>0$ controls the concentration.
For our initial sampling tests for the curved vMF target, we consider $\mathbb{S}^{2}$ and choose $\kappa=500$. Again, we run all four MCMC samplers. Rejection-based geoSSS and HMC take roughly the same amount of time, with HMC taking slightly longer. However, both are almost twice as slow as shrinkage-based geoSSS. RWMH is by far the fastest method (refer to Appendix I for run times).
Figure 8: First panel from left shows the curved vMF target density on the $2$-sphere with concentration parameter $\kappa=500$. Remaining panels show the last 5000 samples from the four different MCMC samplers targeting this density.
Figure 8 shows the target density and samples of the four MCMC algorithms generated in the last 5000 iterations. The underlying target density shown here is evaluated by uniformly distributing $3\times 10^{4}$ points on the $2$-sphere using the method of Saff1997. Visual inspection shows that the slice samplers and also HMC explore the target much more rapidly than RWMH, which is also reflected in the high autocorrelation (see Fig. 9A). To verify this observation, we measure the discrepancy between the discretized target and the histogram generated by the sampling procedure using KL divergence (see Fig. 9B). Moreover, we monitor the distance between successive samples (see Fig. 9C). Across all three evaluation we consider, RWMH is clearly disfavored. On the other hand, geoSSS and HMC show a similar performance on this low-dimensional target.
@@
Figure 9: (A) ACF of the first dimension plotted for the curved vMF target on $2$-sphere based on $10^{5}$ samples. (B) KL divergence between a discretized version of the curved vMF target on 2-sphere and a histogram compiled from the MCMC samples. (C) Geodesic distance (log-scale) between successive approximate samples for the curved von Mises-Fisher distribution on $2$-sphere (17).
However, this picture changes as the dimensionality of the target increases from $d=3$ to $d=5$. We compute the relative ESS for the first dimension by varying the concentration parameter. By using $10$ chains, fixing the dimension $d=5$ and drawing $10^{6}$ samples per chain, we observe (see Fig. 10A), that HMC consistently outperforms the remaining samplers across all values of $\kappa$. The trend of decreasing ESS with increasing $\kappa$ aligns with results from the previous section on the mixture of vMF distributions. HMC is expected to perform exceptionally well on such targets, even if the dimensionality is increased further. To assess the performance of the remaining samplers under similar conditions, we repeat the ESS calculations, however in this case, we fix the concentration parameter with $\kappa=500$ and vary the dimensions from $3$ to $24$. As observed in Fig. 10B, while the ESS values of the slice samplers are better than RWMH, all three exhibit rapid decline with increasing dimension. As expected, in contrast, HMC demonstrates superior performance, maintaining fairly consistent ESS values across all dimensions.
Figure 10: Comparing the relative ESS values. (A) Fixing the dimension $d=5$ and varying the concentration paramater. (B) Fixing $\kappa=500$ and varying the dimensions.
@@ 5 Summary
We introduce two slice sampling based MCMC-methods on the sphere, the ideal geodesic and geodesic shrinkage slice samplers, that use movements on great circles. For the Markov kernels of both samplers we are able to establish reversibility and explicit convergence statements under mild assumptions. In numerical experiments, we see that in particular for moderately concentrated target distributions with a moderate number of dimensions our slice samplers perform well and can compete with RWMH and HMC on the sphere, while having the additional advantage of being free of tuning parameters. For multimodal target distributions we observe that the ideal geodesic slice sampler and the geodesic shrinkage slice sampler even outperform RWMH and HMC. However, we observe that the performance of our methods deteriorates for increasing concentration and increasing dimension of the target distribution. This dependence on concentration and dimension already appears in Theorem 6 and Theorem 16. Regarding the comparison between the two geodesic slice samplers, the ideal version seems to outperform the shrinkage based sampler in terms of Markov chain transitions. However, this comes with a (significantly) greater computational cost per transition.
@@
We have implemented our geoSSS algorithms, including the spherical RWMH and HMC methods (detailed in Appendices F and G), in our geosss package available at https://github.com/microscopic-image-analysis/geosss. You can find the scripts used to generate the numerical illustrations in the scripts directory, and the protein example data in the data directory. Also the precomputed results are archived on Zenodo at https://doi.org/10.5281/zenodo.15610462.
@@ Appendix A Simulation of $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}$
Implementing the slice sampling algorithms proposed in this paper involves sampling from $\mathcal{U}_{{\mathbb{S}^{d-2}_{x}}}=\frac{1}{\omega_{d-2}}\,\mu_{x}$ for $x\in\mathbb{S}^{d-1}$. A scheme for performing this is provided in Algorithm 5 and justified as follows. It exploits that $\mathbb{S}^{d-2}_{x}$} can be interpreted “as tilted $\mathbb{S}^{d-2}$, such that it lies in the (d1)(d-1)-dimensional hyperplane orthogonal to xx”. We formalize the arguments and require further objects. For $x\in\mathbb{S}^{d-1}$ let
@@
be the (d1)(d-1)-dimensional hyperplane with normal vector $x$. Since this is a (d1)(d-1)-dimensional linear subspace of $\mathbb{R}^{d}$, we can choose an orthonormal basis $e_{1}^{(x)},\ldots,e_{d-1}^{(x)}\in E_{x}$ of $E_{x}$. The isometric mapping
@@
describes the rotation of $\mathbb{R}^{d-1}$ into $E_{x}$. Therefore it maps the unit sphere in $\mathbb{R}^{d-1}$ to the unit sphere in $E_{x}$, i.e., $\varphi_{x}(\mathbb{S}^{d-2})={\mathbb{S}^{d-2}_{x}}$. Isometries preserve volume, such that
@@
We prove the validity of Algorithm 5 based upon this relationship between $\sigma_{d-2}$ and $\mu_{x}$. For $x\in\mathbb{S}^{d-1}$ we establish an expression for the mapping $y\mapsto y-(x^{T}\!y)\,x$ in terms of $\varphi_{x}$ that describes the projection of $\mathbb{R}^{d}$ onto $E_{x}$. Then, we deduce that Algorithm 5 realizes a transformed normalized standard normally distributed random variable under the isometry $\varphi_{x}$. By (19) this implies the desired result formalized as follows.
@@
$y-({x}^{T}\!{y})\,x$ $=\sum_{i=1}^{d-1}\alpha_{i}e_{i}^{(x)}=\varphi_{x}\left((\alpha_{1},\dots,\alpha_{d-1})^{T}\right)=\varphi_{x}(U^{T}y).$